Determine the average function for each of the following total functions:
a. Total Revenue = 100Q-Q^2
b. Total cost = 1,000 + 10Q + 0.01Q^2
c. Total Profit = 50Q - 0.1Q^2 - 1,000
1
Expert's answer
2018-03-06T09:41:08-0500
a. Total Revenue = 100Q-Q^2
Average Revenue = (100Q-Q^2)/Q=100-Q
b. Total cost = 1,000 + 10Q + 0.01Q^2
Average cost = (1,000 + 10Q + 0.01Q^2)/Q = 1,000/Q + 10 +0.01Q
c. Total Profit = 50Q - 0.1Q^2 - 1,000
Average Profit = (50Q - 0.1Q^2 - 1,000)/Q = 50 – 0.1Q-1,000/Q